OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -7x + 5y = -56 -21x – 2y = -168

OpenStudy (anonymous):

-7x + 5y = -56 -------- (i) multiply by -3 and add to (ii) -21x – 2y = -168 -------- (ii) 21x -15y = 168 and we get -17y = 0 so y = 0/-17 => y = 0 substitute y = 0 in (i) -7x + 5(0) = -56 -7x = -56 x = -56/-7 x = 8 so the unique solution is (8, 0)